OpenStudy (anonymous):

Describe the end behavior of the function f(x)=7x^4-2x+19 by finding lim f(x) x --> infinity and x--> - infinity

OpenStudy (comm.dan):

So you need help with function notation? @krlg11

OpenStudy (anonymous):

f(x)\[7x ^{4}-2x+19\]

OpenStudy (anonymous):

I need help with where to start! I am not even sure how to tackle this!

OpenStudy (comm.dan):

K I will help u @krlg11

jimthompson5910 (jim_thompson5910):

In general, if the degree of the polynomial function f(x) is odd, then \[\Large \lim_{x\to\infty} f(x) = \infty\] and \[\Large \lim_{x\to -\infty} f(x) = -\infty\]

jimthompson5910 (jim_thompson5910):

and if the degree of f(x) is even (f(x) is some polynomial), then \[\Large \lim_{x\to\infty} f(x) = \infty\] and \[\Large \lim_{x\to -\infty} f(x) = \infty\]

jimthompson5910 (jim_thompson5910):

so you first need to identify the degree of \[\Large 7x^{4}-2x+19\]

OpenStudy (comm.dan):

Do they give u the x or do you have to find the x? @krlg11

OpenStudy (comm.dan):

The degree is 4, which is quartic @jim_thompson5910

jimthompson5910 (jim_thompson5910):

the degree is 4 like Comm.Dan is saying so the degree is even, which means you use this rule \[\Large \lim_{x\to\infty} f(x) = \infty\] and \[\Large \lim_{x\to -\infty} f(x) = \infty\]

OpenStudy (anonymous):

So the degree comes from the first part of the equation?

OpenStudy (comm.dan):

Yes from the first monomial @krlg11

jimthompson5910 (jim_thompson5910):

basically what this means is that as x gets bigger in the positive direction, then y gets bigger as well also x gets bigger in the negative direction, then y gets bigger as well no matter what, y will always get bigger and head off to infinity

OpenStudy (anonymous):

Hmm, okay and knowing that it will equal -infty is just something you have to know?

jimthompson5910 (jim_thompson5910):

well you can use a graph to help out here is some generic odd degree polynomial